OpenStudy (anonymous):

Solve for x. -x + 4 = 32

OpenStudy (anonymous):

-x +4 =32

OpenStudy (anonymous):

32-4 = 28

OpenStudy (anonymous):

28/-1 = -28

OpenStudy (anonymous):

x = -28

OpenStudy (tkhunny):

Or, -x + 4 = 32 Add x 4 = 32+x Subtract 32 -28 = x

OpenStudy (anonymous):

:P that works...

OpenStudy (tkhunny):

As I have said MANY times, unique answers don't care how you find them. Personally, I prefer moving negative values rather that multiplying by negative values. It makes INequalities so much simpler! Anyway, always keep your eyes open for multiple avenues.

OpenStudy (tkhunny):

I know, I know, some people will panic when they see an x on the right hand side. I suggest such individuals simply get over it. ;-)

OpenStudy (anonymous):

haha :P I've completed algebra 2 so it doesn't bother me. ;)

OpenStudy (tkhunny):

"=" is reflexive! x = -28 is the same as -28 = x Good work.

OpenStudy (anonymous):

Haha math just doesn't click in my head. Ive tried everything but when it comes to math, just looking at the problem scares me

OpenStudy (anonymous):

Don't let it scare you. Tell yourself you love it... That's what I did... I got a 100 in Algebra 2

OpenStudy (tkhunny):

ahahaha I sucked in Algebra II. It was more of a social problem in high school, though. Once I woke up, all was well. You don't want to know what year that was. :-)

OpenStudy (anonymous):

haha

OpenStudy (tkhunny):

Okay, 1972! What were you doing in 1972?!

OpenStudy (anonymous):

Wasn't even born. ;)

OpenStudy (tkhunny):

It's okay. Some of us managed to survive.

OpenStudy (anonymous):

haha